Two equally charged objects of mass 1.68 kg are located 0.08 m apart.  Determine the possible charges on the objects if the net force on each of them is zero.




1
Expert's answer
2021-05-27T08:31:23-0400
"k\frac{q^2}{r^2}=G\frac{m^2}{r^2}\\(9\cdot10^9)q^2=(6.67\cdot10^{-11})1.68^2\\q=1.45\cdot10^{-10}\ C"